My Question
A tin which is 12 cm long, 9 cm wide and 4 cm deep holds 120 g of tea. If 1 kg of the same tea is packed into a tin which has a 12 cm square base, how tall will the tin have to be ?
Please help me to solve this. Thanks.

Answer by venugopalramana(3286)

A tin which is 12 cm long, 9 cm wide and 4 cm deep holds 120 g of tea. If 1 kg of the same tea is packed into a tin which has a 12 cm square base, how tall will the tin have to be ?
VOLUME OF TIN =L*B*H=12*9*4=432 CC
IT HOLDS 120 GM.OF TEA
TO HOLD 1 KG=1000 GMS..WE NEED
1000*432/120 =36000 CC
BASE AREA OF NEW TIN =A=12*12=144 SQ.CM.
HENCE HT NEEDED =V/A=36000/144=250 CM
Please help me to solve this. Thanks.